# Retinomorphic Machine Vision in a Network Laser

### v3

@@ Abstract
With the growing prevalence of AI, demand increases for efficient machine learning hardware. Physical systems are sought which combine image feature detection with the essential nonlinearity for tasks such as image classification. Existing physical hardware typically detects features linearly, then employs digital processing for nonlinear activation. Biological vision systems excel at nonlinear image processing. The retina detects features in ganglion cells via lateral inhibition, where cells nonlinearly compete for neuronal firing activity while suppressing neighbouring cells.
We present a bio-inspired ‘retinomorphic’ machine vision platform using an on-chip semiconductor network laser. The system detects multiple features in parallel via spatially-overlapping lasing modes, with integrated nonlinearity provided by antagonistic gain competition between modes - a photonic analogue of retinal inhibition. Parallel feature-detection enhances efficiency relative to feature-detection schemes which operate sequentially or via multiple device copies, with Si-compatible processing and a compact $\upmu$m-scale footprint relative to existing mm-scale systems. We report 98.05% accuracy on MNIST-digits and 87.85% on Fashion-MNIST, with strong performance on short training datasets.
@@ Introduction
Machine learning and its benefits are increasingly ubiquitous across industry and society. However, they are accompanied by exponentially increasing energy and carbon costs. As conventional CMOS processors struggle to keep pace with the rising demands of machine learning, the need increases for dedicated hardware solutions that implement machine learning architectures.
An architecture with pressing need for such solutions is the convolutional neural network (CNN), widely used in machine vision[1, 2]. Machine vision, where tasks such as feature detection, classification and segmentation are performed on image data, represents a significant portion of modern machine learning demand[3, 4]. CNNs begin with a feature detection layer where convolutional kernel filters are raster-scanned across an image to build ‘feature maps’ of image features such as edges, corners and curves. Each kernel detects a single separate feature, with multiple kernels scanned sequentially to produce a set of feature maps. These feature maps are then processed via nonlinear activation and passed to a subsequent classification layer, typically fully-connected hidden layers including additional nonlinear activations. The initial convolutional feature detection greatly enhances machine vision performance relative to neural networks containing only multi-layer perceptron style hidden layers[1, 2, 3, 4].
While powerful, convolutional feature detection is computationally expensive[5, 6]. As such, developing dedicated physical hardware which implements CNN functionality is an active research effort. Recent breakthroughs have seen impressive implementations of convolutional feature detection in physical systems including memristor arrays[7], photonic integrated circuits[8, 9] and meta-surfaces[10], demonstrating the benefits of physical convolutional processing. These existing schemes adopt the structure of software CNNs - performing feature detection in a linear regime by implementing physical equivalents of the linear matrix multiplications used in software.
The drawback of such linear physical approaches is that they are unable to provide the crucial nonlinearity required for more complex machine vision tasks beyond feature detection, such as image classification. This is typically navigated by passing the detected feature maps to a digital CMOS computer which implements nonlinear activation in software[11, 12, 13]. Often, the software also performs the full classification process via a conventional deep neural network with additional nonlinear activations[8, 10]. A multilayer optical scheme has recently been demonstrated using saturating opto-electronic nonlinearity between two linear optical layers, employing a fully-connected rather than convolutional architecture[14].
Biological vision systems function quite differently to software CNNs. Biological feature detection directly integrates nonlinear processing rather than implementing it as a separate subsequent step, and multiple features are detected in parallel with nonlinear coupling between the detection processes[15, 16] as opposed to the discrete kernel filters used in software CNNs. In the retina, feature detection is performed by retinal ganglion cells, the final output neurons in the retinal circuitry[15]. These cells utilise a process termed ‘lateral inhibition’ to nonlinearly enhance image features such as edges[17, 18, 19]. When a ganglion cell is activated, it sends excitatory signals to the brain while simultaneously sending inhibitory neurotransmitter signals to adjacent ganglion cells, decreasing the firing activity of neighbouring cells. This dynamic interplay between adjacent cells provides nonlinear enhancement of image edges, as the inhibition disproportionately affects cells receiving weaker stimuli (e.g., darker image regions). At the boundary between adjacent light and dark regions (e.g., an image edge), brightly illuminated ganglion cells strongly inhibit their weakly illuminated neighbours while receiving less inhibition themselves, thereby nonlinearly amplifying neuronal activity at image edges. Additionally, different subsets of ganglion cells respond preferentially to different angles and orientations of edges, allowing the retina to process complex images with multiple features in parallel[15, 16].
It is appealing to draw inspiration from biological architectures, and engineer physical platforms capable of detecting multiple features in parallel with integrated nonlinearity. Neuromorphic computing is an approach which takes this bio-inspired angle and aims to implement machine learning by leveraging the complex nonlinear dynamics of physical systems[20, 11, 13]. However, gaining fine control over the nonlinear physical dynamics powering neuromorphic computing is a subtle and challenging task, particularly for more sophisticated multi-layer learning architectures such as CNNs where each layer of the neuromorphic system must perform very different processing[13, 12]. To date, a physical system integrating convolutional feature detection and physical nonlinearity remains unrealised to the best of our knowledge. Gaining enhanced control over nonlinear physical dynamics to realise such functional integration has been identified as a key goal for the field[11, 12, 13].
In photonic systems, complex lasers which host many competing lasing modes[21, 22, 23] offer an inviting candidate system for implementing bio-mimetic machine vision. They provide nonlinear processing via lasing threshold dynamics and mode competition[24, 25, 26, 27], and provide spatial sensitivity through their responsive reaction to structured light.[24, 27, 25, 28]. Crucially, nonlinear mode competition for gain between overlapping modes provides an attractive photonic analogue of retinal lateral inhibition processes - offering routes to feature detection. These promising systems have not yet been explored for computation.
Here, we demonstrate a neuromorphic machine vision platform using a lithographically-patterned InP network laser[29], fabricated through the mature process of wafer-bonding InP on oxide-coated Si[30]. InP is a well-established photonic platform suitable for active devices in the telecoms range and robust to degradation. By harnessing nonlinear mode competition, we engineer a bio-inspired ‘retinomorphic’ design where spatially-distributed lasing modes overlap throughout the network, competing for gain and suppressing neighbouring modes in a similar manner to lateral inhibition in retinal ganglion cells. Different lasing modes win the competition process when specific image features are present, providing feature detection by increasing their lasing amplitude - akin to the increased neuronal activity providing retinal feature detection. We integrate convolutional feature detection and strong physical nonlinearity in a single compact 150 $\upmu$m physical system, with ten image features detected in parallel by different lasing modes.
For image classification, detected feature maps are then re-illuminated onto the network to provide a subsequent classification layer with further physical nonlinear processing. A single logistic regression step is applied to the spectral output of this second layer, linearly multiplying each spectral channel by a single fixed weight to provide classification results. Accuracy scores of 98.05% and 87.85% are reported for MNIST digits and Fashion MNIST image classification tasks respectively. Strong classification performance is observed where the system is trained on small datasets, attractive for in-the-field learning applications such as edge-computing. The photonic system outperforms multiple software neural network architectures including CNNs and multilayer perceptrons for training sets below 5,000 (MNIST digits) and 10,000 (Fashion-MNIST) images. Close correlation is found between experimental results and theoretical modelling, highlighting nonlinear mode-competition between overlapping modes as the underlying physical mechanism responsible for feature detection.
@@
Figure 1: Image feature detection in a complex network laser. a) Digital micromirror device spatially structures 633 nm pump light into an arbitrary input image, which is illuminated onto a 150 $\upmu$m InP network b). The pump light triggers optical gain in illuminated regions of InP network, exciting photons which scatter through the network to produce many lasing modes at different wavelengths determined by the scattering paths. Spatial mode profiles of three modes are indicated via different coloured nodes. c) Lasing spectra emitted by the network are recorded via spatially-resolved spectrometer, giving ‘hyperspectral’ data with spectral and spatial dimensions , at $\sim$1 $\upmu$m spatial resolution. Spectral regions which detect image features are highlighted by coloured boxes, corresponding to feature maps in i) and j). d) Mode intensity vs. pump power for all modes, with different thresholds and nonlinear behaviours seen between modes. Colour-coded modes correspond to the colour-coded spectral regions and feature maps in i,j). e) Schematic of feature detection via kernel filters in software convolutional neural networks. Kernel filters are raster-scanned over the input image in small ’windows’ performing sequential multiply-and-accumulate operations to produce feature maps. Each image feature (e.g. left or right edges) is detected by a separate kernel filter, detecting multiple features consumes more time and energy. The process is linear, so subsequent nonlinear activations must be performed on the feature maps to enable tasks such as image classification. f) Schematic of network laser feature detection. Small input image windows are sequentially illuminated on the InP network. Different network lasing modes detect multiple image features in parallel from a single illumination, with integrated physical nonlinearity. Experimentally, ten features are concurrently detected with forty detected in simulation. The network laser efficiently integrates multiple steps of the feature detection process, which must be performed sequentially in software. g) Input ‘cameraman’ image, inset shows example image windows (blue boxes) and raster-scan direction. h) Image windows from g) with spectral output of the InP network and mode intensities for six feature-detecting lasing modes (colours follow spectral regions and feature maps in c,i,j). Mode amplitudes vary in response to the presence of different image features. i) Composite feature map detected by the InP network. Different feature colours correspond to the lasing amplitude of the six lasing modes indicated in c,h,j). A range of image features are detected in parallel. j) Feature maps for the six separate lasing modes which form the composite map, highlighting the range of parallel features detected by the system.
@@
Figure 1e) shows a schematic of feature detection in conventional software CNNs. A kernel filter is raster scanned across small image ‘windows’ of an input image, performing multiply-and-accumulate operations via matrix multiplication. The kernel values are multipled against the image pixel values, the results summed to a single number which becomes the pixel value of an output feature map at the centre of the image window location. This process is repeated across the input image to build feature maps, with different kernel filters sensitive to different image features e.g. left or right edges. To detect multiple image features, the process is repeated sequentially using different kernels - consuming additional time and energy for each feature detected. This process is entirely linear, so subsequent nonlinear activations must be performed at time and energy cost to use the feature maps for ML tasks such as image classification. This linear approach is what has so-far been used for physical convolutional feature detection schemes[8, 9, 7, 10].
Figure 1f) shows a schematic of the network laser scheme. Image windows are illuminated on the network in a raster-scan fashion similar to CNNs, but here the network simultaneously detects many features in parallel via multiple lasing modes, with integrated physical nonlinearity. This increases efficiency relative to software CNNs and their physical counterparts, with the network laser currently detecting ten parallel features experimentally (forty in simulation) in a compact 150 $\upmu$m footprint, relative to the mm-scale of alternative hardware schemes.
Figure 1g-j) show experimental parallel feature detection for a ‘cameraman’ test image (g). Example image windows ($\times 4$ pixels, inset of g) are shown in (h) with their respective network lasing spectra and mode intensities for six feature-detecting modes. Mode lasing intensity is increased or suppressed by the presence of different image features due to a combination of correlation between the input image shape and spatial mode profile, and nonlinear mode competition between spatially-overlapping modes. Specific modes lase strongly when a given image feature is present (e.g. a left or right hand edge), providing direct feature detection via the lasing amplitude. This enables feature maps to be obtained by plotting the lasing amplitude of different modes at different positions in the raster-scan. Figure 1i) shows a composite feature map produced by six lasing modes, with the different colours corresponding to the lasing amplitude of the six modes indicated in a,e,g). A broad range of features are detected in parallel, with separate output for each feature provided by the different spatial and spectral locations of the modes. Figure 1j) shows separate feature maps for the six modes comprising the composite map. Feature maps for all ten feature-detecting modes are shown in supplementary figure S10.
@@
a) 50 $\upmu$m network topology used for netSALT simulations a-f). b) Three network illumination patterns: uniform, horizontal edge and vertical edge. c) Simulated network lasing spectra for the three illumination patterns. Three modes A-C are indicated, corresponding to d). d) Simulated lasing mode intensities of three modes A-C for the three network illumination patterns. Mode competition nonlinearly suppresses different modes below lasing threshold and amplifies other modes dependent on input image. Spatial mode profiles are shown inset in the left panel. e) Input image, ‘CL’ for ‘complex lasing’. The input image is fractured into 8$\times$8 pixel windows and illuminated onto the network sequentially in a raster-scan fashion, similar to kernel scanning in a convolutional filter process. f) Simulated feature maps for five modes. Pixel brightness in the feature maps corresponds to the mode lasing intensity for that position in the window raster-scan. Different lasing modes respond to different image features, spatial mode profiles are shown inset. A larger set of 28 simulated feature maps corresponding mode profiles is shown in the supplementary information. g) Experimentally recorded ‘CL’ feature maps produced from six modes. As in simulation, different lasing modes respond to different image features. h) Experimental composite feature map comprising modes from g). i) Equivalent 4$\times$4 kernels for the six lasing modes, extracted by fitting experimental data. Coloured frames correspond to mode feature map colour scheme. $R^{2}$ scores suggest the effective linear kernels fail to capture some of the nonlinear physical dynamics. j) Photographic test image of a street scene, 300 $\times$ 200 pixel resolution. input image. k) Experimental composite feature map of the street test image.
Spectrally-multiplexed feature-detection
@@
We can use physical modelling to examine the origin and functionality of the photonic feature-detection. Here, we employ a numerical model netSALT[25] which solves the nonlinear interaction of optical waves on graph networks within the steady-state ab-initio laser theory (SALT)[31] approximation. The netSALT model includes amplification and loss on graph edges and nonlinear competition between spatially-overlapping lasing modes. Figure 2a) shows a random Voronoi graph network used for the netSALT simulations. Network topology and average edge length (5 $\upmu$m) match the experimentally measured network shown in fig. 1, with a smaller 50 $\upmu$m diameter used in simulation to reduce computational load.
@@
Three illumination patterns are explored: uniform illumination (‘full’), horizontal edge (‘right’) and vertical edge (‘top’) (fig. 2b). The resultant simulated netSALT lasing spectra are shown for the three illuminations in figure 2c), showing strong variation in lasing spectra under the different illumination patterns.
@@
NetSALT predicts 239 possible lasing modes for the network considered here, with 172 attaining lasing threshold for at least one of the illumination patterns. The three spectra in fig. 2c) show the sum of all active modes for the different illumination patterns. To gain a better understanding of the behaviour of specific modes which drive the feature-detection process, we examine three specific modes, labelled A,B,C, identified in figure 2c) by dashed vertical lines. Figure 2d) shows the lasing amplitude of the three modes under ‘full’, ‘right’ and ‘top’ illumination patterns with the spatial mode profiles inset in the left-hand panel. The modes are spatially-distributed across the network, with non-zero spatial overlap between the three modes. As multiple modes are active in the same regions of the network, nonlinear mode-competition occurs between the modes as they antagonistically compete for gain. Under full illumination, mode B shows the highest amplitude, while mode C does not lase at all as it falls below the lasing threshold. For a right-edge illumination mode B now falls below the lasing threshold and is deactivated, allowing mode C to lase with the highest amplitude as it no longer faces strong mode-competition from the mode B. For a top-edge illumination, mode A exhibits far higher amplitude and mode C now falls below the lasing threshold.
@@
It is important to note that modes A and C exhibit higher lasing amplitude under ‘top’ and ‘right’ edge illuminations, where the network receives less input pump energy than under uniform illumination. This is not possible in a linear system with non-interacting modes, and occurs here as removing pump light from some network regions causes modes residing even partially in those regions to lose the mode-competition interactions, e.g. mode B winning under full illumination and losing under edge illumination. This observed behaviour shows that nonlinear competition for gain between spatially-overlapping modes is providing feature-sensitive lasing dynamics, analogous to the lateral inhibition processes in retinal ganglion cells.
We now consider the netSALT-simulated feature-detection performance of the system. An input image of the letters CL (complex lasing) is shown in fig. 2e). The image was fractured into 8x8 pixel windows and illuminated on the network in a raster-scan fashion. The lasing response was simulated for each window illumination, and the mode lasing amplitude plotted at each position in the raster scan to generate separate feature-maps for each mode.
@@
Figure 2f) shows feature maps provided by five different modes, with the corresponding spatial mode profiles inset. These plots demonstrate the ability of the network to perform spectrally-multiplexed feature-detection of a range of image features, with the broad set of lasing modes effectively providing parallel convolutional filters. The spatial location of the mode profiles within the network are seen to show correspondence with the detected image features, e.g. modes detecting lower image edges are located towards the bottom of the network, modes detecting left edges are located towards the left of the network, and so on. Of the 172 active modes in the simulated network, 40 were identified as exhibiting good feature-detection. A set of 28 selected feature maps and corresponding spatial mode profiles is shown in the supplementary figure S3, alongside the full set of 172 feature maps from all modes in supplementary figure S4. A regular hexagonal network was simulated to compare against the random graph topology (supplementary figure S5), with far fewer detected features in the regular network showing the benefits of a disordered graph topology. This demonstrates the ability of network topology to control the device functionality, with further explorations into the correspondence between topology and performance fruitful grounds for future work.

### v4

# Few-Shot Retinomorphic Vision in a Nonlinear Photonic Network Laser

@@ Abstract
With the growing prevalence of AI, demand increases for hardware that mimics the brain’s ability to extract structure from limited data. In the retina, ganglion cells detect features from sparse inputs via lateral inhibition, where neurons antagonistically suppress activity of neighbouring cells. Biological neurons exhibit diverse heterogenous nonlinear responses, linked to robust learning and strong performance in low-data regimes.
Here, we introduce a bio-inspired ‘retinomorphic’ photonic system where spatially-competing lasing modes in a network laser act as heterogeneous, inhibitively-coupled neurons - enabling few-shot classification and segmentation. This compact (150 $\upmu$m) silicon-compatible scheme addresses key challenges in photonic computing: physical nonlinearity and spatial footprint. We report 98.05% and 87.85% accuracy on MNIST and Fashion-MNIST, and 90.12% on BreaKHis cancer diagnosis - outperforming software CNNs including EfficientNet in few-shot and class-imbalanced regimes. We demonstrate combined segmentation and classification on the HAM10k skin lesion dataset, achieving DICE and Jaccard scores of 84.49% and 74.80%. These results establish a new class of nonlinear photonic hardware for versatile, data-efficient neuromorphic computing.
@@ Introduction
Artificial intelligence (AI) is increasingly ubiquitous across society and industry. Demand grows for dedicated AI hardware capable of strong performance in use cases where training data and energy are scarce, such as edge computing applications where limited data is collected locally and the power and cooling overheads of GPUs are unattractive. These tasks, including image classification, medical diagnosis, and segmentation, require a range of capabilities:
  • Strong nonlinearity – crucial for harder tasks[1, 2, 3].
  • Extraction of a range of informative features from input data – neural networks capable of this such as convolutional neural networks (CNNs) [4, 5] and transformers[6, 7] typically outperform those without, e.g. multilayer perceptrons (MLPs) [4, 5, 8, 9].
  • The ability to rapidly learn from few training examples – termed the few-shot/low-data regime [10, 11].
Physics-based neuromorphic computing systems[1, 12, 3, 13, 14, 15, 16] have been explored as candidates for edge-AI hardware[14], but to date demonstrations integrating the above qualities and providing software-equivalent performance on harder problems such as biomedical tasks have remained elusive.
Biological systems offer a compelling blueprint. In the retina, ganglion cells perform spatially-sensitive nonlinear processing via lateral inhibition[17, 18, 19, 20], where neurons compete to fire while inhibitively suppressing activity of neighbouring cells. This mechanism enables efficient encoding of spatial structure, even under sparse or noisy input, and the ability to detect many features in parallel from multiple ganglion cells[17, 21]. The combination of excitatory dynamics - firing in response to input, and inhibitory ones - suppressing activity of nearby coupled neurons, dynamics is crucial to the strong performance of biological learning systems, but to date physics-based neuromorphic schemes typically focus solely on engineering excitatory dynamics. Addressing this is an inviting route towards taking neuromorphic performance closer to the biological systems which inspire it. In addition, the diverse heterogenous nonlinear responses across ensembles of biological neurons have been linked to robust learning[22], and creating effective high-dimensional latent spaces which can give rise to strong few-shot learning[11].
Among photonic systems, complex random lasers hosting many interacting modes[23, 24, 25, 26] offer promising neuromorphic platforms. Their lasing thresholds, multiple diverse modes, and competition for optical gain between modes[27, 26, 28, 29] provide intrinsic heterogenous nonlinearity, missing or weak in many photonic schemes[3, 30, 13, 2], while their sensitivity to structured optical input provides spatial processing[27, 29, 26, 31]. The mode competition inhibitively couples modes, as they antagonistically try to harvest a finite amount of pump energy. Despite these properties, such systems have remained unexplored for computation.
Here, we present a retinomorphic machine vision system implemented in a 150 $\upmu$m diameter on-chip semiconductor random network laser[32], fabricated through the mature process of wafer-bonding InP on oxide-coated Si, followed by electron beam lithography and reactive ion etching[33]. The large number of spatially-overlapping lasing modes provide effective heterogenous photonic neurons, performing parallel image feature detection through their spatially-sensitive response, with nonlinearity arising from competition for optical gain between neighbouring coupled modes - a photonic analogue of retinal inhibition. Beyond feature extraction, we demonstrate higher-level vision tasks of classification and segmentation, with strong performance in few-shot/low-data regimes. Crucially, the system outperforms a range of software neural networks including EfficientNet, a large-scale modern CNN, on hard biomedical tasks with scarce and class-imbalanced training data.
These results position the system not just as a novel photonic device addressing the size and nonlinearity issues facing optical computing, but as an exemplar of a new class of compact, few-shot capable neuromorphic hardware. By embedding bio-inspired behaviour combining excitation, inhibition, and nonlinear neuronal heterogeneity directly into the physical dynamics, we enable neuromorphic processing that meets the demands of edge AI.
@@
Figure 1: Working principle of neuromorphic network laser vision scheme. a) Digital micromirror device spatially structures 633 nm, 200 fs pulse pump light into an arbitrary input image, which illuminates a 150 $\upmu$m InP network b). The pump light triggers optical gain in illuminated network regions, exciting photons which scatter through the network to produce many lasing modes at different wavelengths determined by the scattering paths. Spatial profiles of three lasing modes are indicated via different colour network nodes. c) Lasing spectra emitted by the network are recorded via spatially-resolved spectrometer, giving ‘hyperspectral’ data with spectral and spatial dimensions, at $\sim$1 $\upmu$m spatial resolution. Spectral regions which detect image features are highlighted by coloured boxes, corresponding to feature maps in i) and j). d) Schematic of feature detection via kernel filters in software convolutional neural networks. Linear kernel filters are raster-scanned over the input image in small ‘windows’, performing sequential multiply-and-accumulate operations to produce feature maps. Each image feature (e.g. left or right edge) is detected by a separate kernel filter sequentially. As the process is linear, subsequent nonlinear activations must be performed on feature maps to enable vision tasks such as image classification. e) Schematic of network laser feature detection. Small windows of input image are illuminated on the network in a raster-scan. Different lasing modes detect image features in parallel, with integrated physical nonlinearity. Experimentally, ten features are concurrently detected. f) Input ‘Cameraman’ image, inset shows example 4$\times$4 image windows (blue boxes) and raster-scan direction. g) Image windows from f) with corresponding network spectral output (2nd col.), LL plots showing the heterogenous nonlinear response of network modes in response to the different input image pump illuminations (3rd col.), and histograms showing mode intensities for six feature-detecting lasing modes (4th col.). Mode amplitudes vary in response to the presence of different image features, enabling feature detection. h) Composite feature map detected by the InP network. Different feature colours correspond to the lasing amplitude of the six lasing modes. i) Feature maps from the six lasing modes which form the composite map (h), highlighting the range of parallel features detected.
@@
Figure 1d) shows a schematic of feature detection in conventional software CNNs. A kernel filter is raster-scanned across small image ‘windows’ of an input image. The kernel values are matrix multiplied against image pixel values, and the results summed to a single number which becomes the pixel value of an output feature map, centered at the image window location. This process is repeated across the input image to build feature maps, with different kernel filters sensitive to different image features e.g. left or right edges. To detect multiple image features, the process is repeated sequentially using different kernels - consuming time and energy. This process is entirely linear, with subsequent nonlinear activations performed at time and energy cost in order to use the feature maps for higher-level AI tasks such as image classification or segmentation. This linear approach is what has so-far been employed in physical convolutional feature detection schemes[34, 35, 36, 37].
Figure 1e) shows a schematic of the network laser feature detection. Image windows are illuminated on the network in a raster-scan fashion similar to CNNs, but here the network simultaneously detects many features in parallel via multiple lasing modes, with integrated physical nonlinearity. One scan produces many feature maps in parallel, ten experimentally (forty in simulation). Our network occupies a compact 150 $\upmu$m footprint, relative to the mm/cm-scale of alternative physical feature-detecting schemes.
Figure 1f-i) show experimental parallel feature detection results for a ‘Cameraman’ test image (f). Example image windows ($\times 4$ pixels, inset of f) are shown in (g) alongside their respective lasing hyperspectra (2nd col.); LL plots (3rd col., network lasing mode intensity vs pump intensity) which show strongly heterogenous nonlinear responses for each mode, and a variation in these nonlinearities depending on the input image; and a histogram of mode intensities (4th col.) for six feature-detecting modes, taken at the pump power indicated by dashed vertical lines in the LL plots. Mode lasing intensity is increased or suppressed by the presence of different image features, due to a combination of correlation between the input image shape and spatial mode profile, and nonlinear mode competition between spatially-overlapping modes. This enables feature maps to be obtained by plotting the lasing amplitude of different modes above the uniform illumination level at different positions in the raster-scan. The mode intensities plotted in the histogram in (g) are used to generate the pixel amplitudes of our feature maps. Figure 1h) shows a composite feature map produced by six lasing modes, with different colours corresponding to the lasing amplitude of six modes indicated in fig. 1c,g-i). Figure 1j) shows separate feature maps for the modes comprising the composite map. Feature maps for all ten feature-detecting modes are shown in supplementary information.
@@
a) 50 $\upmu$m network topology used for netSALT simulations a-f). b) Three pump illumination patterns: uniform, horizontal edge (‘right’) and vertical edge (‘top’). c) Simulated network lasing spectra for the three illumination patterns. Three modes A-C are indicated, corresponding to d). d) Simulated lasing intensities of three modes A-C, for the three pump illumination patterns. Mode competition nonlinearly suppresses specific modes below lasing threshold, and amplifies other modes dependent on input pump image. Spatial mode profiles are shown inset in left panel. e) Input image, ‘CL’ for ‘Complex Lasing’. The input image is fractured into 8$\times$8 pixel windows and illuminated onto the network sequentially in a raster-scan fashion, similar to kernel scanning in a convolutional filter process. f) Simulated feature maps for five modes. Pixel brightness in the feature maps corresponds to mode lasing intensity for that position in the window raster-scan. Different modes lase in response to different image features, spatial mode profiles shown inset. A larger set of 28 simulated feature maps and corresponding mode profiles is shown in supplementary information. g) Experimentally measured ‘CL’ feature maps produced from six modes, with different lasing modes responding to specific image features. h) Experimental composite feature map combining modes from g). i) Equivalent 4$\times$4 kernels for the six lasing modes in (g,h), extracted by linear fits to experimental data. Coloured frames correspond to mode feature map colour scheme. j,k) Colour photograph of a mandrill. Image is separated into RGB channels and feature detection is performed at multiple kernel sizes demonstrating the resolution flexibility of our scheme. Composite RGB feature maps are shown in (k) for 4$\times$4, 7$\times$7 and 11$\times$11 kernels illustrating the range of high-to-low spatial frequency features detected as kernel size is increased.
Parallel Feature-Detection via Network Lasing Modes
@@
We can use simulations of the photonic physics to examine the origin of the feature-detection. Here, we employ netSALT[26], a numerical model which solves the nonlinear interaction of optical waves on graph networks within the steady-state ab-initio laser theory (SALT)[38] approximation. The netSALT model includes amplification and loss on graph edges and nonlinear competition between spatially-overlapping lasing modes. Figure 2a) shows a random Voronoi graph network used for the netSALT simulations. Network topology and average edge length (5 $\upmu$m) match the experimentally measured network shown in fig. 1, with a smaller 50 $\upmu$m diameter used in simulation to reduce computational load.
@@
Three illumination patterns are explored (fig. 2b): uniform illumination (‘full’), horizontal edge (‘right’) and vertical edge (‘top’). The resultant simulated lasing spectra are shown in figure 2c), with strong variation in lasing spectra under the different illumination patterns.
@@
NetSALT predicts 239 possible lasing modes for the network considered here, with 172 reaching lasing threshold for at least one of the illumination patterns. The three spectra in fig. 2c) show the sum of all active modes for the different illumination patterns. We focus on three specific modes, labelled A,B,C, identified in figure 2c) by dashed vertical lines. Figure 2d) shows the lasing amplitude of the three modes under ‘full’, ‘right’ and ‘top’ illumination patterns with the spatial mode profiles inset in the left-hand panel. The modes are spatially-distributed across the network and overlap considerably. As multiple modes are active in the same network waveguides, nonlinear mode-competition occurs between the modes as they antagonistically compete for finite pump gain. Under full illumination, mode B shows the highest amplitude, while mode C does not lase as it falls below threshold. For a right-edge illumination, mode B now falls below the lasing threshold and is deactivated, allowing mode C to lase with the highest amplitude as it no longer faces strong mode-competition from the mode B. For a top-edge illumination, mode A exhibits far higher amplitude and mode C now falls below the lasing threshold.
@@
It is important to note that modes A and C exhibit higher lasing amplitude under ‘top’ and ‘right’ edge illuminations, where the network receives less input pump energy than under uniform illumination. This is not possible in a linear system with non-interacting modes, and occurs here as removing pump light from some network regions causes modes residing even partially in those regions to lose the mode-competition interactions, e.g. mode B wins under full illumination and losing under edge illumination. This observed behaviour shows that nonlinear competition for gain between spatially-overlapping modes is providing feature-sensitive lasing dynamics, analogous to the lateral inhibition processes in retinal ganglion cells.
@@
We now consider the netSALT-simulated feature-detection performance of the system. An input image of the letters CL (‘Complex Lasing’) is shown in fig. 2e). The image was fractured into 8x8 pixel windows and illuminated on the network in a raster-scan fashion. The lasing response was simulated for each window illumination, and the mode lasing amplitude plotted at each position in the raster scan to generate separate feature-maps for each mode. Figure 2f) shows feature maps provided by five different modes, with the corresponding spatial mode profiles inset. These plots demonstrate the ability of the network to perform spectrally-multiplexed feature-detection of a range of image features, with the broad set of lasing modes effectively providing parallel convolutional filters. The spatial location of the mode profiles within the network are seen to show correspondence with the detected image features, e.g. modes detecting lower image edges are located towards the bottom of the network (2nd panel), modes detecting left edges are located towards the left of the network (4th panel), and so on. Of the 172 active modes in the simulated network, 40 were identified as exhibiting good feature-detection. A set of 28 selected feature maps and corresponding spatial mode profiles is shown in the supplementary information, alongside the full set of 172 feature maps from all modes in supplementary information. A regular hexagonal network was simulated to compare against the random topology (supplementary information), with far fewer detected features in the hexagonal network. This demonstrates the ability of network topology to control the device functionality, with further explorations into the correspondence between topology and performance providing fruitful ground for future work.
Figure 3: Network Laser Image classification. a) Single-layer classification. Input images are illuminated directly onto the network, hyperspectra are measured, and a single logistic regression step is applied to the spectra to provide classification. Accuracy of 96.03% is observed for MNIST-digits.
b,c) Multi-layer classification. An initial feature detection layer (b) is implemented as described in figure 2, generating 10 feature maps in parallel via the amplitude of ten lasing modes (indicated by coloured boxes on the hyperspectra). A second classification layer (c) is then implemented by illuminating the feature maps sequentially and measuring hyperspectra. The hyperspectra are combined, and a single logistic regression step applied to return classification predictions. Accuracy of 98.05% is observed for MNIST-digits.
d,e,f) Three classification tasks are implemented, with example images for each task showing detected features: (d) MNIST-digits, (e) Fashion-MNIST, and (f) BreaKHis cancer diagnosis (400x magnification).
g,h,i) Classification accuracy vs. training set length, comparing single- and multi-layer photonic network accuracy against software benchmarks: logistic regression, multilayer perceptron (MLP), LeNet-5 convolutional neural network (CNN), a larger CNN (96 filters), and the far larger EfficientNetV2-B0 CNN (3040 filters, in a transfer-learning scheme with pre-trained ImageNet weights). Maximum accuracies for the multi-layer photonic network of Digits: 98.05%, Fashion: 87.85%, and BreaKHis: 90.12% are observed. Zoomed-in regions of the few-shot regime are shown in (g,h), where the photonic network exceeds software benchmark performance.